OpenStudy (anonymous):

I really need major help with word problems... Please if anyone can guide me with how these things are worked out, would really appreciate: The first problem : Fred drove home to the airport 320 km away at an average speed. If he had increased his average speed by 1km/h, he would have taken 20 mins less. What was his average speed? This is under the topic of quadratic equation word problems. Please help. Thanks

OpenStudy (ajprincess):

Let us say the average speed is s. The distance between the home and the airport is 320km. Also let us say that time taken to cover that distance with the speed s is t. speed=distance/time s=320/t t=320/s the increased speed=s+1 new time=t-20 distance is the same. s+1=320/(t-20) (s+1)*(t-20)=320 t-20=320/(s+1) t=320/(s+1)+20 320/(s+1)+20=320/s does this help?

OpenStudy (ajprincess):

@miszery

OpenStudy (anonymous):

okay thanks let me study it a bit lol i'm slow

OpenStudy (ajprincess):

oh k. take your own time:)

OpenStudy (anonymous):

ooh okay i see cool :)

OpenStudy (anonymous):

do you want to help me with more word problems @ajprincess :PP

OpenStudy (ajprincess):

well sure:) have u solved this one?

OpenStudy (anonymous):

30.5 km /h ?

OpenStudy (anonymous):

trying to figure out how to use the quadratic equation to solve that atm

OpenStudy (ajprincess):

pls wait a sec

OpenStudy (ajprincess):

ooops reallly sorrry. I made a mistake. let me correct it

OpenStudy (anonymous):

yeah teh teacher has only the answer but not showing the working :(

OpenStudy (anonymous):

okay

OpenStudy (ajprincess):

Let us say the average speed is s. The distance between the home and the airport is 320km. Also let us say that time taken to cover that distance with the speed s is t. speed=distance/time s=320/t t=320/s the increased speed=s+1 upto this 5n.

OpenStudy (anonymous):

here is the answer she gave: 320/x - 320/x+1 = 1/3 how did she get 1/3 ?

OpenStudy (ajprincess):

20 minutes will be 1/3 of a hour. 60minutes in an hour. 20/60=1/3

OpenStudy (anonymous):

ooh

OpenStudy (anonymous):

righto thats clearer L:D

OpenStudy (anonymous):

can u show me how to turn this into a quadratic equation?

OpenStudy (anonymous):

\[320x + 320 - 320x \div x ^{2}+x = 1/3 (x ^{2}+x) \right?\]

OpenStudy (anonymous):

ew yuk

OpenStudy (anonymous):



OpenStudy (anonymous):

oh ok

OpenStudy (ajprincess):

is that clear?

OpenStudy (anonymous):

yep thanks :)))

OpenStudy (ajprincess):

using quadratic formula u can solve it. do u knw it?

OpenStudy (anonymous):

yes i know this part hehehe.. just putting the numbers into that equation is hard for me

OpenStudy (anonymous):

but once its in that form I can solve it

OpenStudy (ajprincess):

so the new time will be t-1/3. distance is the same. s+1=320/(t-20) (s+1)*(t-20)=320 t-1/3=320/(s+1) t=320/(s+1)+1/3 320/(s+1)+1/3=320/s 1/3=320/s-320/(s+1) =320(1/s-1/(s+1)) =320((s+1-s)/s(s+1)) =320(1/(s(s+1))) 1*(s(s+1))=320*3 s^2+s=960 s^2+s-960=0

OpenStudy (ajprincess):

reallly sorrry for messing it up. nw is the above one clear? it is the correct one. if it is nt cleat pls ask.

OpenStudy (ajprincess):

Did ur teacher give u the final answer for the average speed?

OpenStudy (ajprincess):

@miszery

OpenStudy (anonymous):

yes sorry 30.5 km

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

hmm. i guess that's the answer u will get u solve the quadratic equation s^2+s-960=0